# Darkness cannot bind them: a no-bound theorem for $d=5$ Myers-Perry null & timelike geodesics

### v2

@@ III.2 Radial motion
$C=$ $4\mu^{2}\left(b\Phi+a\Psi\right)^{2}+4\left(a^{2}-b^{2}\right)\left[\left(x+b^{2}\right)\Phi^{2}-\left(x-a^{2}\right)\Psi^{2}\right]-4\Delta\left[m^{2}\left(x+\frac{a^{2}+b^{2}}{2}\right)+Q\right]\,.$
@@
Physical motion requires $\mathcal{X}\geq 0$, therefore, since $A>0$, the energy, $E$, must satisfy either $E<V_{-}^{{\rm eff}}$ or $E>V_{+}^{{\rm eff}}$. In principle these two regimes should be considered separately, however performing the substitutions
@@
one effectively transforms $V_{\pm}^{{\rm eff}}\rightarrow V_{\mp}^{{\rm eff}}$, thus one can consider only $E>V_{+}^{{\rm eff}}$ and then extend the analysis to $E<V_{-}^{{\rm eff}}$ by means of the above transformation.
@@
In order to have radially bounded motion one must have at least three radial turning points, i.e. three distinct roots of the function $\mathcal{X}$, outside the event horizon for any value of $E>0$, since $V_{+}^{{\rm eff}}\rightarrow 0^{+}$ at infinity. This reasoning is illustrated in Fig. (1), where a dummy potential is depicted with three turning points for some values of $E$.
@@
Notice that possibility $(ii)$ is impossible, since if $c_{2}<0$, then $c_{0}>0$, as it is a sum of all positive terms 30. Now possibility $(ii)$ will be considered, which implies that bounded motion requires $E^{2}<m^{2}$, as it happens in $4$ dimensional spacetimes.
@@
Now to assert the impossibility of having radially bound timelike geodesics in the $d=5$ MP spacetime it will be proven that it is impossible to simultaneously satisfy $c_{1}<0$ and $\mathcal{P}>0$. Since $\mathcal{P}$ is a convex parabola and $c_{1}$ can only be negative between −$-\alpha p/\mu^{2}$ and −$-\beta q/\mu^{2}$, then $\mathcal{P}$ must be positive at at least one of these points, otherwise it is negative in the entire interval violating the necessary condition for radially bound motion. At $E=-\alpha p/\mu^{2}$ one has

### v3

@@ III.2 Radial motion
$C=$ $4\mu^{2}\left(b\Phi+a\Psi\right)^{2}+4\left(a^{2}-b^{2}\right)\left[\left(x+b^{2}\right)\Phi^{2}-\left(x+a^{2}\right)\Psi^{2}\right]-4\Delta\left[m^{2}\left(x+\frac{a^{2}+b^{2}}{2}\right)+Q\right]\,.$
@@
Physical motion along the radial direction requires $\mathcal{X}\geq 0$, therefore, since $A>0$, the energy, $E$, must satisfy either $E\leq V_{-}^{{\rm eff}}$ or $E\geq V_{+}^{{\rm eff}}$. However, these conditions alone do not fully determine whether a trajectory is physically admissible. By analogy with four-dimensional spacetimes, one expects that physically meaningful trajectories should satisfy $E\geq 0$ at spatial infinity. It is important to emphasize that the conditions $E\leq V_{-}^{{\rm eff}}$ and $E\geq V_{+}^{{\rm eff}}$ are independent and do not necessarily imply $E\geq 0$ or $E\leq 0$, respectively. Depending on the choice of parameters, it is possible to have regions where $V_{+}^{\mathrm{eff}}<0$ or $V_{+}^{\mathrm{eff}}>0$. Meaning that these conditions are independent and must both be considered for a full analysis of radial motion. However performing the substitutions
@@
one effectively transforms $V_{\pm}^{{\rm eff}}\rightarrow-V_{\mp}^{{\rm eff}}$, thus one can consider only $E\geq V_{+}^{{\rm eff}}$ and then extend the analysis to $E\leq V_{-}^{{\rm eff}}$ by means of the above transformation.
@@
In order to have radially bounded motion one must have at least three radial turning points, i.e. three distinct roots of the function $\mathcal{X}$, outside the event horizon for any value of $E>0$, since $V_{+}^{{\rm eff}}\rightarrow m^{+}$ at infinity. This reasoning is illustrated in Fig. (1), where a dummy potential is depicted with three turning points for some values of $E$.
@@
Notice that possibility $(i)$ is impossible, since if $c_{2}<0$, then $c_{0}>0$, as it is a sum of all positive terms 30. Now possibility $(ii)$ will be considered, which implies that bounded motion requires $E^{2}<m^{2}$, as it happens in $4$ dimensional spacetimes.
@@
Now to assert the impossibility of having radially bound timelike geodesics in the $d=5$ MP spacetime it will be proven that it is impossible to simultaneously satisfy $c_{1}<0$ and $\mathcal{P}>0$. Since $\mathcal{P}$ is a convex parabola and $c_{1}$ can only be negative for energy $E$ between −$-\alpha p/\mu^{2}$ and −$-\beta q/\mu^{2}$, then $\mathcal{P}$ must be positive at at least one of these points, otherwise it is negative in the entire interval violating the necessary condition for radially bound motion. At $E=-\alpha p/\mu^{2}$ one has
@@ VI Conclusions and discussion
Another interesting extension of our analysis is the case of charged rotating black holes. It is easy to verify that the addition of charge does alter the possibility of radially bound geodesics in static five-dimensional black hole spacetimes. However, whether charge could qualitatively change the behaviour in the spinning case remains an open question. Unfortunately, no fully analytical solution for a charged rotating Myers-Perry black hole in five dimensions is known. While a perturbative slowly spinning solution exists [29], it does not admit full separability of the geodesic equations and lacks the Carter-like constant, implying that the motion is not integrable. As a result, our analytical approach cannot be directly applied to this case. Addressing this question would require numerical methods to analyse geodesic motion in such spacetimes, which lies beyond the scope of this work but is an interesting avenue for future investigation.